OpenStudy (anonymous):

Evaluate: [(21 + 6) − 3^2] ÷ 9 • 2= is the answer 4

OpenStudy (anonymous):

If you do: \[21+6=27\]then: \[27-9=18\]then:\[18/9=2\]then:\[2*2=4\]

OpenStudy (anonymous):

so im rite

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Use PEMDAS Parentheses Exponents Multiply Divide Add Subtract. So yes you are correct

OpenStudy (anonymous):

yay thank you all

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

I call it BIDMAS: Brackets Indices Division Multiplication Addition Subtraction But the same idea.

OpenStudy (anonymous):

depends on where you are :)

OpenStudy (skullpatrol):

Summary of Order of Operations: 1. First simplify expressions within grouping symbols. 2. Then simplify powers. 3. Then do all multiplications and divisions in order from left to right. 4. Then do all additions and subtractions in order from left to right.